Listed below are speeds (in mi/h) of cars on the New Jersey

Turnpike by race of the driver. All cars are going in the same direction, and all of the cars are from New Jersey. The data are from Statlib and the authors are Joseph Kadane and John Lam berth. Find Variance . and find the standard deviation.

79    70    71 76 76    74    71 75 74 74

Group of answer choices

a) s=2.749, Var = 7.55

b) Var= 74 s=2.749

c) mean = 7.4 s=2.749

d) mean = 74 s=2.749
Answer #1

The sample varince is given by

s2 = (ir - T)/(n -1)

where

} - 740/10 = 74

Thus , variance is

s2 2 =Σ(α – 7402/(10 – 1) = 68/9 = 7.55

Standard deviation is

s= V7.555 = 2.749

Answer is

a) s= 2.749 , Var = 7.55

Calculation :

x (x-74)^2
79 25
70 16
71 9
76 4
76 4
74 0
71 9
75 1
74 0
74 0
sum 740 68